TERM HOWTO

Patrick Reijnen patrickr@bart.nl v1.2, 15 July 1995

Contents

1 Legal Information 3 1.1 Copyright statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 DISCLAIMER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Introduction 3 2.1 This Document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 What is TERM? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 How TERM works 4 3.1 Nomenclature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 Setting Things Up 5 4.1 What has to be available . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4.2 Explanation of concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4.2.1 Sharing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4.2.2 Full TERM networking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4.3 Build TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.3.1 Build TERM, versions 2.0.0 and higher . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.3.2 Build TERM, versions 1.16 up to 1.19 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.3.3 Build TERM up to version 1.15 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

4.4 client.a, libtermnet.a, libtermnet.sa, libtermnet.so . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.5 Setting environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.6 Test TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.7 TERM and communication programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

4.7.1 Kermit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.7.2 Seyon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.8 Make a Transparent Link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.9 Run linecheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.10 Try Running TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.11 Terminate your TERM connection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.12 Removing TERM from your partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.13 Optimizing your connection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.14 Troubleshooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

5 TERM clients 21

6 X and TERM 22

7 tredir  23

7.1 tredir can bite! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.2 Stupid tredir tricks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

7.2.1 X windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.2.2 Mail with TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

8 tudpredir  29

9 Automating Things 29

10 Porting software for use with TERM 31 10.1 Port and compile the sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 10.2 Termify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

11 Term clients 33 11.1 Term clients available on ftp sites. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 11.2 The termnet package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 11.3 Asked for but not yet supported . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

12 Term and Security 35 12.1 trsh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 12.2 txconn and xauth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 12.3 sxpc, xhost and xauth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 13 Things to remember 37

14 Reliability of TERM versions 38

15 Term speed table 38

16 Hints and Tricks found on the net 40

17 Other Things 41

18 Acknowledgements 41

1 Legal Information

1.1 Copyright statement

This document may be distributed freely as a whole in any form and free of charge. Parts of this document may be distributed, provided that this copyright message is included and the reader is informed that this is not the full HOWTO document. Furthermore, there is to be a pointer as to where the full document can be obtained. Specifically, it may be included in commercial distributions, without prior consent. However, I would like to be informed of such usage.

This HOWTO may be translated into any language, whatsoever, provided that you leave this copyright statement and the disclaimer intact, and that a notice is appended stating who translated the document.

1.2 DISCLAIMER

While I have tried to include the most correct and up-to-date information available, I cannot guarantee that usage of the information in this document does not result in loss of data. I provide NO WARRANTY about the information in this HOWTO and I cannot be made liable for any consequences for any damage resulting from using information in this HOWTO.

2 Introduction

2.1 This Document

This HOWTO attempts to clear up some of the confusion of using TERM, Michael O’Reilly’s remarkable program that allows you to multiplex your serial line and set up a network connection. By and large, the documents that come with TERM are quite good, and this HOWTO is not intended to replace them. The intention of this document is to give some background on how TERM works and detail the steps in getting some of the more common networking services working under TERM. It should be emphasized that this document does not cover everything there is to know about TERM. After reading it, the TERM manual pages should be read, since they include information not contained here.

2.2 What is TERM?

TERM is a program, written by Michael O’Reilly (michael@iinet.com.au) and maintained by Bill Riemers (bcr@physics.purdue.edu), that is run over a serial line to allow multiple connections to operate concurrently - i.e. you may be down-load a file via your modem while working on a (different) remote system via the same modem connection. TERM can also be used to open up X client windows over a serial connection. Through the tredir utility and the tudpredir utility TERM can provide almost all of the “traditional” TCP/IP and UDP network services: mail, news, ftp, telnet, xarchie, etc. In a sense, TERM is very much like other serial protocols such as SLIP or PPP. TERM’s advantage is that it can be run entirely from user space, requiring no support from system or network administrators.

Unlike SLIP or PPP, your machine does not have its own IP address. All incoming traffic must be addressed to your remote host, and it will be redirected to your local computer by TERM.

3 How TERM works

Before experimenting with TERM it is strongly advised to first read this complete chapter and the INSTAL-LATION file provided with the package. Also take a look at the manual pages linecheck, (term)test and TERM. This will help you to work easier and faster.

3.1 Nomenclature

I assume you are dialling a system through some sort of terminal server. I use the terms local and remote to refer to the home and network connected systems respectively (unless I use them to mean something else :-).

TERM provides the local machine, which has no network connection, but is connected, via a serial line, to a remote machine which is in turn connected to a network with network services. Let us look at how a machine with a ’traditional’ network connection provides these services. First the user invokes a program, like ftp or telnet, that requests a network service. What these programs do is make a system call requesting network services. The operating system then obtains these services via its network interface (e.g. it sends and receives packets over the ethernet). SLIP and PPP do exactly this, by converting your modem line into a network interface, which is in principle no different from an ethernet. The downside of this is that these protocols make the modem-connected machine part of the network, just like any other machine. This implies all the administrative burdens associated with being a network node (more actually, since the modem link must also be administered).

In the absence of a network connection like SLIP or PPP, what does one typically do? Well, you dial your network connected machine, read your mail, your news etc; if you need a file, you first transfer it to the remote machine and then download it to your local machine using kermit or some other communication program. This is a bit of a pain, especially since you can only really do one thing at a time that uses your modem link. The idea behind TERM is basically to automate and multiplex this process. TERM is invoked on both the local and remote machines, and the two processes communicate with one another over the modem line. When you need a network service, you make a request to the local TERM daemon, which forwards the request to the TERM daemon on the remote, network-connected, machine. The result is then returned over the modem line.

To be more concrete, say you want to retrieve a file by ftp. First you need a version of ftp that can speak to TERM. You invoke this termftp as you do a regular ftp, say termftp nethost.gov, but this special version makes its network request to the local TERM daemon instead of the kernel. The local TERM forwards this request, over the modem line, to the remote TERM, which opens an ftp connection to nethost.gov, and transmits the data back over the modem link. TERM is smart enough to have many different things going on at once; so you can have several different network sessions using the same modem link, e.g. you can be logged into another distant host via termtelnet while the termftp transfer is going on.

If this is too abstract (or unclear) do not worry; the important piece of information to get out of this section is that there are two copies of TERM running, one on each end of the modem link.

4 Setting Things Up

4.1 What has to be available

Before you start building and using TERM you have to make sure that you have built TCP/IP support into the kernel. Furthermore, make sure that the TCP/IP loopback interface is activated. When this is the case you can go on with the rest of this section.

4.2 Explanation of concepts

In newer TERM version two new concepts have entered TERM. These two concepts will be explained in the next two subsections.

4.2.1 Sharing

Starting with version 1.16 the concept of sharing the TERM connection with other users has entered TERM. This means that when you enable shared features, other people can use the same TERM connection you are using, i.e. when you are working on your remote machine via your TERM connection (say, from your local machine you used trsh to get in) another person on your local machine can use the same TERM connection at the same time to ftp a file to his login on your local machine from an ftp site somewhere in the world.

When you disable shared features (i.e. you execute TERM in private mode) you and only you (we do not count root :-) can use the TERM connection.

Of course, you only need to install shared TERM at the end at which you want to allow people to use the same TERM connection you are using. So, if other people have a login on your local machine and they want to use it from somewhere on your remote network you enable shared features on the remote end of your TERM connection. In this way all these people can login on your machine at the same time sharing the same TERM connection with each other and with you. (NOTE: the first example needed shared features to be enabled at the local end of the TERM connection).

NOTE for installation as root: When you install TERM as root you have to create a ’term’ group first (before compilation) with no member by adding the following line in ’/etc/group’:

  term::16:root

or any other unused GID than 16 when 16 is already in use.

After compilation and installation make TERM and its clients SGID ’term’:

  chgrp term <term_client>
  chmod g+s <term_client>

Also any other program that you make TERM-aware must be made SGID ’term’.

4.2.2 Full TERM networking

Starting with TERM version 2.0.0 the statement full TERM networking is used. When your only connection with the outside world is a TERM connection, you have a full TERM network and you should build TERM with full TERM networking. In this case in the shared directory a file called termnet is placed. This tells TERM that your only connection to the outside world is via TERM.

When you also have some other type of network connection beside your TERM connection TERM-aware programs first try to fulfil their job using this network connection. When this fails TERM is invoked and it is tried to fulfil the job via the TERM connection. To make this more clear now an example is given in which TERM-aware telnet is used. This telnet should work both with and without TERM.

  telnet localhost

does not use TERM to connect, but

  telnet zeus.cs.kun.nl

will use TERM only if you do not have some other type of network connection.

Full TERM networking also means to lie about the host name, and say it is the remote host instead. Furthermore, it causes bind(0) to always act on the remote host. In essence it makes many programs unusable when they are not going through TERM, while TERM is running. Unfortunately, most UDP programs and daemons will not work with TERM without these nasty tricks.

4.3 Build TERM

When you are lucky, this should just involve a make. Most probably however, you need to do more. Due to new features in newer versions of TERM it is now a bit more complicated to create your TERM binary. A couple of ways can be followed to obtain your binary.

To cover all these ways TERM can be built this section will be split into three parts:

1. Build TERM, versions 2.0.0 and higher

2. Build TERM, versions 1.16 up to 1.19

3. Build TERM up to version 1.15

4.3.1 Build TERM, versions 2.0.0 and higher

First, make sure you have read the section about ’full TERM networking’ above.

For TERM versions 2.0.0 and higher there are many ways to build the TERM binary and the clients. All of these can be done both by root and by ordinary user:

1. Build TERM in private mode without full TERM networking

2. Build TERM in private mode with full TERM networking

3. Build TERM in shared mode without full TERM networking

4. Build TERM in shared mode with full TERM networking

In these versions of TERM a new way for compilation has entered TERM using the script configure. When configure is run it checks on what operating system you are trying to install TERM, whether the source directory is available or not, and if any runtime options are set. According to the things found configure then creates a Makefile using Makefile.in which is provided in the TERM package.

Two of the more important options to configure are --root and --user which state whether TERM will be installed by root or an ordinary user. Other options can be used to install TERM the way you want (non-standard paths for example).

1. Build TERM in private mode without full TERM networking. To build TERM in this way you need to execute the following commands (both for root and ordinary user):

       ./configure --root  OR --user
       make install installman

This builds the binaries and installs these binaries and the manual pages.

2. Build TERM in private mode with full TERM networking. To build TERM in this way you need to execute the following commands (both for root and ordinary user):

       ./configure --root  OR --user
       make installnet installman

This builds the binaries and installs these binaries and the manual pages.

3. Build TERM in shared mode without full TERM networking.

To build TERM in this way you need to execute the following commands (both for root and ordinary user):

  ./configure --root  OR --user
  make share installman

This builds the binaries and installs these binaries and the manual pages.

4. Build TERM in shared mode with full TERM networking.

To build TERM in this way you need to execute the following commands (both for root and ordinary user):

  ./configure --root  OR --user
  make share installnet installman

This builds the binaries and installs these binaries and the manual pages.

4.3.2 Build TERM, versions 1.16 up to 1.19

To build these versions of TERM you can now choose one of the following ways:

1. As an ordinary user, build TERM in private mode

2. As an ordinary user, build TERM in shared mode

3. As root, build TERM in private mode

4. As root, build TERM in shared mode

Below, it will be explained how to enable/disable shared features during the compilation of TERM.

1. You are an ordinary user (no root access) and you do NOT want to SHARE the TERM connection with other users. As a user who does not want to share the TERM connection with other users you should do the following to build TERM:

  make DO=install OS-type
  make installman

After this TERM, its clients and the manual pages are built and installed. Furthermore, you need to create a directory ’$HOME/.term’. This is the directory in which TERM will look for its ’termrc’ file.

2. You are an ordinary user (no root access) and you want to SHARE the TERM connection with other users. As a user who wants to share the TERM connection you should do the following:

  make DO=installshare USERSHARE=$HOME/term OS-type
  make installman

After this TERM, its clients and the manual pages are built and installed. Furthermore, you will have a directory ’$HOME/term’ (default) with permissions ’drwxrwxr-x’. In this directory you will find at least the socket used by TERM for its connection (’tmp/private/socket=’).

3. You are root and you do NOT want to SHARE the TERM connection with other users. As root who does not want the TERM connection to be shareable you should do the following to build

TERM:
  make DO=install OS-type
  make installman

After this TERM, its clients and the manual pages are built and installed. Furthermore, you now have a directory called ’/usr/local/lib/term’ (default) with permissions ’drwxr-xr-x’. In this directory you will at least find the socket used by TERM for its connection (’tmp/private/socket=’).

4. You are root and want to SHARE the TERM connection.

First, make sure you have read the section about ’sharing’ above. As root who wants to share the TERM connection you should do the following:

  make DO=installshare OS-type
  make installman

After this TERM, its clients and the manual pages are built and installed. Furthermore, you now have a directory called ’/usr/local/lib/term’ (default) owned by group TERM and with permissions ’drwxrwxr-x’. In this directory you will at least find the socket used by TERM for its connection (’tmp/private/socket=’).

4.3.3 Build TERM up to version 1.15

For these versions of TERM building should invoke no more than the commands

make DO=install OS-type
make installman

You will find TERM, its clients and the manual pages nicely built and installed and ready for use after this.

Furthermore, you need to create a directory ’$HOME/term’. This directory TERM will use to look for its termrc file.

The only thing you may want to do is change some of the paths in the Makefile or change some of the compiler flags.

4.4 client.a, libtermnet.a, libtermnet.sa, libtermnet.so

With TERM a library with functions for TERM clients is provided.

Up to version 1.16 this library was called client.a. During compilation of TERM this library was built and then used during the compilation of the TERM clients. It was not installed in another directory.

Starting with version 1.16 the name of the library is changed to libtermnet.a. Up to version 1.19 this library is created in the TERM directory and then used during compilation of the TERM clients. It is not installed in another directory.

Starting with version 2.0.0, beside libtermnet.a also libtermnet.so and libtermnet.sa (shared library and exported initialized library data) are created during compilation of the TERM package. During the installation of all the parts of the package also these three library files are installed in the directory ’/usr/local/lib’ (default). Then a link is made from libtermnet.so.2 to libtermnet.so.2.x.x. Finally, ldconfig is run to create the necessary links and cache (for use by the run-time linker, ld.so) to the most recent shared libraries found in the directories specified on the command line, in the file ’/etc/ld.so.conf’, and in the trusted directories (’/usr/lib’ and ’/lib’). If the installation is done correctly the three library files can now be used by TERM clients which are built with dynamic instead of static libraries. Also, these libraries can now be used to port your own software to make it TERM aware (see belo! w).

4.5 Setting environment variables

TERM knows a couple of environment variables which can be set by users. The first three of those that I will explain are

• TERMDIR

• TERMSHARE

• TERMMODE

By setting these variables you can control the way TERM is run.

For TERM versions up to 1.15 only the variable TERMDIR is important (these versions do not know the shared mode). For these versions TERMDIR should be set as follows:

  setenv TERMDIR $HOME     (csh or tcsh)
  export TERMDIR=$HOME     (bash)

Starting with version 1.16 TERM also knows the variables TERMSHARE and TERMMODE. With these variables TERM can be told to run in private mode or in shared mode. I will explain how to set the variables for private mode and shared mode.

TERMMODE knows the following three values;

• 0 = private

• 1 = system shared

• 2 = user shared

1. Running TERM in private mode can be done setting the variables TERMDIR and TERMMODE in the following way: For csh or tcsh

       setenv TERMDIR $HOME
       setenv TERMMODE 0

For bash

       export TERMDIR=$HOME
       export TERMMODE=0

2. When you want to use TERM in shared mode there are two ways of setting the variables:

(a) When TERM is installed as a SUID program only TERMMODE has to be set.

           setenv TERMMODE 2    (csh or tcsh)
           export TERMMODE=2    (bash)

(b) When TERM is installed as a SGID program the variables have to be set in the following way: For csh or tcsh

           setenv TERMMODE 1
           setenv TERMDIR /usr/local/lib/term
           setenv TERMSHARE $TERMDIR

For bash

           export TERMMODE=1
           export TERMDIR=/usr/local/lib/term
           export TERMSHARE=$TERMDIR

Setting the variables in this way makes it possible to start old clients (clients linked to an older version of client.a) in shared mode.

Starting with version 2.0.0 TERM also knows the variable TERMSERVER. You need to set this variable when you have multiple modems and you have more than one connection at a time. To specify which connection to use, you must start TERM with a server name:

  nohup term -v /dev/modem1 Connection1 &
  nohup term -v /dev/modem2 Connection2 &

Users should then set the variable TERMSERVER to the connection name they want to use:

  setenv TERMSERVER Connection1   (csh or tcsh)
  export TERMSERVER=Connection2   (bash)

4.6 Test TERM

Do a make test (or make termtest for newer versions of TERM) to build TERM’s test daemon. (term)test works by running two copies of TERM on your system, a ’local’ and a ’remote” copy. Both of these will read your ’termrc’; so you can adjust their behaviour. Now execute (term)test. You should now be able to do a

trsh and a tupload (try
  tupload ./term /usr/tmp

- you should get a copy of the TERM binary in ’/usr/tmp’). The local TERM’s output should show up in ’local.log’, the remote’s one in ’remote.log’. You can start TERM up with a -d255 flag to enable debugging output to be written to these files, or enable debugging in your ’termrc’ file.

NOTE: Run test as ./test so as to avoid your system’s test.

4.7 TERM and communication programs

Before you can use TERM you must have established a connection via the modem using a communication program like kermit or seyon. In the documentation of your communication program you can find what you need to do to establish the conection with the remote machine.

when you have established the connection with the remote machine and you want to run TERM you need to suspend or quit your communication program without closing the connection with the remote machine. This needs to be done as otherwise the communication program will steal characters from linecheck or TERM.

Below for some communication programs I will explain how you can make sure that the connection will stay alive and the communication programs will not steal characters from linechech or TERM.

4.7.1 Kermit

Starting TERM when you use kermit is easy. At the local kermit prompt you type suspend. Now you see back your Linux prompt. From this prompt you can establish your TERM conection.

4.7.2 Seyon

An easy way to start linecheck or TERM when you are using seyon is to put linecheck and TERM in the Transfer Menu (controlled by the file ’$HOME/.seyon/protocols’).

In the file ’$HOME/.seyon/protocols’ add:

  "Line check" "$cd /tmp; linecheck"
  "Term" "$term -c off -w 10 -t 150 -s 38400 -l $HOME/tlog"

Then, when you want to execute linecheck or TERM on the local machine, you can select the Transfer Menu, either the ”Line check” or the ”Term” item, and Go.

Of course, you can also use the shell command button, and type ’linecheck’ or ’term’ in the pop-up dialog box. This also does automatic redirection of input and output.

4.8 Make a Transparent Link

Presumably, you can establish a modem connection between your local and remote hosts. Typically, you are dialling into some kind of terminal server and connecting to your remote host from there. You are also using some kind of terminal software, such as kermit or seyon to talk to your modem (the examples in this document will use kermit, since that is what its author uses). If you are having trouble with your modem, or your terminal software, take a look at the Serial-HOWTO; that should help you out.

Having established your link, you want to make it as transparent as possible. Check the commands on the terminal server (help or ? will usually get you started). Go for the 8 bit options whenever possible. This may mean changing the way you log in to a system, e.g. if the server uses rlogin, you may have to use it and give it the -8 flag to make it transparent. Especially watch out for xon/xoff flow control. You do not want that. Try to enable rts/cts (hardware) flow control. You may need to check your modem documentation to learn how to configure it to do 8-bit rts/cts communications.

4.9 Run linecheck

WARNING: In some of the documents the command line options for linecheck are mentioned in an incorrect order. I have checked this and found the order of options mentioned below to be the correct ones.

NOTE: Starting with TERM version 2.3.0 linecheck no longer needs to have the name of a log-file on its command line. It will write its output to the file ’linecheck.log’ in the directory you start linecheck in.

Linecheck is a program that is supplied with TERM. It checks the transparency of a link, providing configuration information that TERM needs to run correctly. linecheck will send each of the 256 possible eight bit characters over the link and verify that each was transmitted successfully. TERM needs to be configured to deal with characters that cannot be transmitted over the link, and linecheck determines what characters these are. You use linecheck after you have established as transparent a modem link as possible. To run linecheck, do the following

1. On the remote system run

       linecheck linecheck.log

2. Escape back to your local system and suspend your communication program (see above)

3. On the local system run

       linecheck linecheck.log > /dev/modem < /dev/modem

When linecheck is done, you will find a set of numbers at the bottom of the ’linecheck.log’ files. These should be escaped in the termrc at the other end of the link. For example, in my system my local ’linecheck.log’ said nothing and my remote ’linecheck.log’ said to escape 29 and 157. Therefore, my local ’termrc’ escapes these characters and my remote ’termrc’ escapes nothing. If I escape a character at one end, I have to also ignore it at the other; so, in this example, I shall have to ignore 29 and 157 in my remote system.

If linecheck hangs, try using

  linecheck linecheck.log 17 19

on the remote system and

  linecheck linecheck.log 17 19 > /dev/modem < /dev/modem

on the local system. This will escape your xon/xoff (flow control) characters, which will hang your line if you have got software flow control. If this solves the hanging problems, you will want to escape /ignore 17/19 in both ’termrc’s’. If your terminal server has other characters that will shut it down, try running linecheck with those characters escaped, as above. You can spot these characters if linecheck hangs. If this is the case, kill it, then look in the log-files. The last characters transmitted are likely to be the culprits. Try it again with these characters escaped.

In summary, my local termrc has the lines

  escape 29
  escape 157

and my remote termrc has the lines

  ignore 29
  ignore 157

since my remote ’linecheck.log’ said to escape 29 and 157.

4.10 Try Running TERM

Log into the remote system, making the link as transparent as possible (if you have not already done so). Fire up TERM at the remote end. I use the following:

  exec term -r -l $HOME/tlog -s 38400 -c off -w 10 -t 150.

Let us run down each option one by one (note that I could just as easily have put these options in my termrc. I did it this way because it saves editing a file while getting TERM set up).

exec means to destroy your current shell, running the given program in its place. I exec things because I do not intend to use my login shell again; so it is just wasting memory. If you are debugging the link and can reliably kill the remote TERM, you might not want to do an exec.

The -r option is needed at exactly one end. TERM will then see this end as the remote end of the connection (Note that TERM’s remote end can thus be your local machine). If you do not use this option at one end TERM clients will spontaneously crash.

-l $HOME/tlog. This logs errors to the file tlog in my home directory. Very useful for debugging. No reason not to do this.

-s 38400 : I have got a 14400 baud modem, with compression. For optimal compression ratios, I want to be able to push bits down the pipe as fast as possible. For a slower modem, you should use something lower. Note that if you have a slower machine with 16450 uart on your serial port, high baud rates can cause data loss by overloading the chip on your serial port. TERM will recover from this, but if you see a lot of error messages in your log file, (or get overrun warnings from linux kernel versions 0.99pl15 and up) you again might want to lower this number.

-c off : This turns data compression off. I have got a compressing modem, and I do not want to compress things twice.

-w 10 -t 150 : Again, these are options to optimize my fast modem link. I have set my window to 10 and my timeout to 150. This is according to the recommendation in the term setup man page.

Escape back to your local machine and suspend your communication program (see above). You do not want it running while TERM is running, because it will fight with TERM over the serial port. If you can convince your modem to not hang up when you exit your communication program (when it toggles DTR), you could just exit the program at this point.

Now run TERM locally. I use:

  term -c off -l $HOME/tlog -s 38400 -w 10 -t 150 {\(<\)} /dev/modem > /dev/modem \&

I need to tell TERM where the modem is; so I point both standard input and output at ’/dev/modem’ (that is what the < and > do). I also run it in the background; so I can use this screen for something else if I want to.

TERM should work now :-). Try a trsh, and see what happens. If you hang, or your link seems slow, take a look at your ’tlog’ at each end. Are you getting timeouts or error messages? If so, then you have configured something incorrectly. Try again (after you have finished reading this :-). Note that the connection will not seem blazingly fast, especially if you are using compression - it will be a little jumpy. The real speed comes in during file transfer and the like.

4.11 Terminate your TERM connection

Most certainly, after you have done a lot of work using TERM, you want to finish your work and bring your TERM connection down. For this to be realized there are four ways:

1. Kill the TERM programs at both sides of the connection. This is the least recommended way of terminating your connection.

2. A better way is to execute the following command locally:

  echo '00000' > /dev/modem

This will nicely terminate your TERM connection. It will work for all version of TERM. Make sure that the sequence of zeros contains at least five zeros.

3. In the termrc of TERM versions 2.0.0 and higher you can now enter a statement called terminate '<some string>'. This sets a string that will cause TERM to exit (’00000’ by default). It must be at least five characters long, to avoid accidently terminations.

4. Starting with version 1.14 there is the program tshutdown (actually for version 1.14 it is available as a patch, for newer versions it is in the package). Executing tshutdown nicely terminates your TERM connection.

4.12 Removing TERM from your partition

Ok, you asked for this. As some of you want to get rid of TERM I here present you the steps to be done in removing TERM. In the process of removing TERM you have to fulfil the following steps:

• Removing directories with their contents. Depending on how you installed TERM, one or more of the following directorieswill exist on your machine:

  $HOME/.term/termrc
  $HOME/.term/termrc.<server>
  $HOME/term/termrc
  $HOME/term/termrc.<server>
  /usr/local/lib/term/termrc
  /usr/local/lib/term/termrc.<server>
  /etc/termrc
  /etc/termrc.<server>

These directories can be removed together with their contents. Use ’/bin/rm -rf’ to get this done.

• The group ’term’. For some of the ways of installation you had to create a group ’term’. Check the file ’/etc/group’ for the ’term’ entry. When it exists you can remove the entry.

• The TERM-package and TERM-aware executables. This is probably the hardest part in removing TERM. For the executables coming with the TERM-package you have to look in the directory ’/usr/local/bin’ or the directory ’$HOME/bin’. With executables you made TERM-aware yourself I cannot help you. You need to know what executables you made TERM-aware in order to know what executables you have to remove. Do not forget configuration, default and other files coming with some of these executables.

• Library files. To remove these you best can execute the following commands:

       cd /
       find . -name libtermnet* -exec /bin/rm {}

This will find and remove the library files.

• Include file. Also for this one the easiest way is to execute the following two commands:

       cd /
       find . -name termnet.h -exec /bin/rm {}

This will remove the include file.

• Manual pages. When you have installed the TERM manual pages you can now find them back in one of the following directories:

       /usr/local/man/man1
       /usr/local/man/cat1
       $HOME/man/man1
       $HOME/man/cat1

At least you have to check for the following manual pages: term, term clients, term setup, tdownload, linecheck, trdate, trdated, termrc, termtest, tmon, tredir, trsh, tshutdown, tudpredir, tupload, txconnand finally tiptest.

• Temporary user directory. This is de directory ’/usr/tmp/private’ and its contents.

After this exercise you can be quite sure that you have removed everything related to TERM.

4.13 Optimizing your connection

Once you have got TERM running, you might want to try to get things optimized. A good way to measure the speed of your link is to run tmon in one window while up/downloading a file in another. Try both (big) text files and compressed files; the plain text should go a factor of two-ish faster than the compressed files. The parameters you want to fiddle with are baud rate (-s), compression (-c), windows (-w), timeout (-t) and retrain (-A).

Watch out with the retrain parameter. With TERM version 1.19 I got a performance decrease of 80% to 90% compared to running TERM without the retrain parameter. It is not clear if this is a bug in TERM version 1.19 and if this problem exists only with TERM version 1.19.

Baudrate: the maximum number of bits per second TERM will try to send over the serial link. TERM will avoid sending characters at a higher data rate than this. The default is to use the speed of your computer’s serial port, but be warned that this may be too high if your modem runs at a lower rate over the phone line. The baud rate option is intended for systems that buffer output to the modem. During setup and tuning it is better to use a small baud rate rather than one which is too large. For high speed links (> 38400), making it unlimited is probably advantageous. This is achieved by using the value ’off’. TERM will then rely solely on your kernel to do flow control.

Compression: you want this on if you do not have a compressing modem. If you do have such a modem, turn compression off, otherwise you will be compressing things twice, which typically increases the amount of data transmitted. Compressing modems are those that use the MNP-5 or V42.bis protocols. Check your modem documentation and the message when your modem connects.

Windows: this is the number of chunks of data, or packets, that TERM will let go over the line before it gets an acknowledgment (or ack) from the remote TERM. For fast modems, increasing this can be a win; for slower links this can overwhelm the remote end.

Timeout: the time TERM will wait for an ack. If you have increased windows, and you are getting timeouts in your log-file, try increasing this.

For 14400/V42.bis, I use -c off -w 10 -t 150. I get around 1700 cps on compressed files and 3500 cps on ASCII files using tupload.

4.14 Troubleshooting

In this section some thoughts are given about what to check when you have problems executing TERM or one of its clients.

• Did you clean up the TERM directory structure? With newer versions of TERM the structure of the directory tree under ’/usr/local/lib/term’ has changed a couple of times. If you are not aware of this, it can cause all kinds of error messages. The best thing to do is to delete the directory tree under ’/usr/local/lib/term’ (save your ’termrc’) and then install your new TERM version. This way, you avoid the struggle with a messed up directory tree.

• Did you remove old sockets? When you update your TERM version remove all the sockets (called ’socket=’) created by TERM. No doing this can cause strange problems. To find out what socket TERM is listening to you can use the ”netstat” program.

• TERM does not compile correct on sunOS 4.1.3? You have configured TERM with ’./configure –user’. During compilation you are getting a assembler error on a unknown ’-k’ flag. The reason of this error is unknown. The solution to this error is to configure TERM with static libraries. So, you have to do ’./configure –user –static’ and then continue with the compilation process the way you normally do. Now TERM should compile correct.

• termtest is presenting you the error: ’Term: failed to connect to term socket ’/root/.term/sockettest”? When termtest runs it expects the executable ’term’ to be in the same directory as termtest. When you do a ’make install’ prior to running termtest, the TERM binary is moved to ’/usr/local/bin’ (or some other bin directory). The workaround for this is to link the binary to the source directory:

       ln -s /usr/local/bin/term /usr/src/term-<version_number>/term

• Are you running the right binary? TERM has been updated quite a lot, and many systems have different versions of the programs floating around. Make sure you are using the right version. Note that this applies to linecheck too. You can use bash’s type -a or the whereis command to find which program is being run. TERM versions after 1.11 should print out their version number when they start up. (Although version 1.14 claims to be 1.12. Sigh.)

• Do you have the right ’termrc’ in the right place? Depending on the version of TERM you are running and the way you installed TERM (being root or user) this file has to be in one of the following directories:

  $HOME/.term/termrc
  $HOME/.term/termrc.<server>
  $HOME/term/termrc
  $HOME/term/termrc.<server>
  /usr/local/lib/term/termrc
  /usr/local/lib/term/termrc.<server>
  /etc/termrc
  /etc/termrc.<server>

Some systems have pre-installed ’termrc’ files; make sure they are gone before you set things up. If you are running things as root, lookout for ’/.term’. TERM creates files (sockets actually) while it is running; so it has its own directory, ’$HOME/.term’, where the file ’termrc’ goes (note, there is no leading dot in ’termrc’!).

• Does TERM find its ’termrc’ file? when you start up TERM at both sides, you should see messages like the one below:

  Term version: 2.2.9
  Reading file:  /usr/local/lib/term/termrc
  Using shared mode.

When the second line is missing TERM cannot find its ’termrc’ file and you know that something is gone wrong during the installation (unless you are not using a ’termrc’ file and enter all the options to the command line :-). Check the place and the permissions of the ’termrc’ file on the site TERM cannot find its ’termrc’ file.

• Do the entries in the ’termrc’ file have the correct syntax and spelling? A known problem is that people who need to escape and ignore several characters enter them like this in the ’termrc’ file:

  escape 1,4,30,255
  ignore 1,4,30,255

Although TERM does NOT like this it will not complain. It only escapes or ignores the first character presented to the escape or ignore statement. The others characters are silently handled as not being there. When you have to escape or ignore several characters you have to put them on separate lines in your ’termrc’ file, each line starting with the keyword escape or ignore. Only when you have to escape or ignore a range of characters you can do it the following way:

  escape 16-19      # escape characters 16, 17, 18, 19
  escape 23         # 23
  escpae 255        # and 255
  ignore 16-19      # ignore characters 16, 17, 18, 19
  ignore 23         # 23
  ignore 255        # and 255

• Is you ’term’ or ’.term’ directory mounted with NFS? If your ’term’ or ’.term’ directory is mounted with NFS you need to set the flag -DTERM NFS DIR in the CFLAGS line of the Makefile. Alas, for the author this flag causes a compile error when TERM 1.19 is compiled on a machine running sunOS 4.*.

• Are all files and directories owned by the correct user and group and do they have the correct permissions? This should be no problems as these permissions are set during the installation phase. However, when you port your own programs to TERM you must be aware of this. Also when you change the mode TERM is working in (i.e. from private mode to shared mode) file and directory ownerships and permissions have to be adapted.

• You are getting the error gethostbyname: <hostname>: Non-authoritative ‘host not found’, or, server failed To solve this you have to check the following things:

1. Is the file ’/etc/hosts’ configured correctly? <hostname> is not the name of your host (old SLS releases and some old and new Slackware releases are shipped with hostname ’darkstar’ for example). Change this in this file. It must at least contain a line like below (the format is described above it):

      # Local Hosts Format:
      #  IP_NUMBER            HOSTNAME        ALIASES
      #
      # Here is the name of your host, first, followed by any aliases
      127.0.0.1       localhost       linuxpc.domain  linuxpc

When your only connection to the outside world is by TERM the line above is the only one that may appear in the file ’/etc/hosts’. Do not put actual addresses of frequently visited destinations in ’/etc/hosts’. TERM does not like these to be in ’/etc/hosts’ and it will hang on them.

2. Are your ’/etc/rc*’ and ’/etc/resolv.conf’ files world readable (chmod ugo+r)?

3. Last, make sure that you have installed the TCP/IP loopback-interface on your machine. You can check for loopback by running the command ’ifconfig’. When you have installed the loopbackinterface you will at least get something like:

      lo        Link encap:Local Loopback
                inet addr:127.0.0.1  Bcast:127.255.255.255  Mask:255.255.255.0
                UP BROADCAST LOOPBACK RUNNING  MTU:2000  Metric:1
                RX packets:0 errors:0 dropped:0 overruns:0
                TX packets:4984 errors:0 dropped:0 overruns:0

on your screen. During boot time you can furthemore check which protocols are in use. On my machine a line like the one below then appears:

           IP Protocols: ICMP, UDP, TCP

For the installation of TCP/IP and the loopback-interface, please read to the NET-HOWTO.

• You are getting all kinds of ’timed out’ messages in your TERM log files? This means that your TERM connection is not optimized. A small number of these messages every now and then is not a problem. These are most certainly due to temporal factors influencing the physical connection between your local and remote hosts. When you get a lot of these messages all the time your connection will slow down considerably. You have to fiddle with the parameters mentioned in the section ’Optimizing your connection’ above. Alas, this part of the installation is a process of trial and error. No hard rules can be given about the numbers to be presented to the various parameters as many factors have an influence on the connection. These factors differ between the connections and even in time.

• Normal FTP with redirected ports does not work for you? Alas, it is a known problem that redirection of the ports needed by FTP (20 and 21) does not give you a working FTP. The only solution is to get a TERM-aware ftp or ncftp version. Alas, also some TERM-aware ftp versions seem not to work.

5 TERM clients

TERM provides several default clients. They include trsh, tmon, tupload, tredir, txconn and in newer versions trdate, trdated. Furthermore, starting with version 2.0.0 tudpredir is available and from version 2.1.0 tdownload is available. This section will deal with trsh, tmon, tupload, tdownload, trdate and trdated. The others each have their own section. No TERM client will work until you have established a TERM link.

tmon is a simple utility to monitor the statistics of your link. It prints a time histogram of characters transmitted and received. It is invoked simply as tmon. Since around version 1.11, tmon has had a bug that causes some information to be garbled (??).

trsh is similar to rsh. Without arguments, it spawns an interactive shell on the remote system (i.e. it logs you in). trsh is one of the primary means of accessing the remote end of the link via TERM. If given an argument, trsh executes that argument as a command on the remote system. For example

  trsh ls

would give you a listing of the files in your home directory on the remote system.

tupload will transfer a file, given as its first argument, from local to remote. By default, the files will be put in the same directory that you invoked TERM from at the other side. To put files in another directory, give their names as a second argument to tupload. For example, if I want to put a copy of the file ’term114.tar.gz’ in ’/usr/tmp’ on the remote system, I would type

  tupload term114.tar.gz /usr/tmp

.

When you use tupload you can use wild cards like in ’tupload a.*’. The shell expands the wild card and tupload is called as ’tupload a.1 a.2 ......’.

tdownload will transfer a file, given as its first argument, from remote to local. By default, the files will be put in the same directory that you invoked TERM from at the local side. To put files in another directory, give their names as a second argument to tdownload. For example, if I want to put a copy of the file ’term114.tar.gz’ in ’/usr/tmp’ on the local system, I would type

  tdownload term114.tar.gz /usr/tmp

.

When you use tdownload you cannot use wild cards like in ’tdownload a.*’. Reason for this is that the remote directory is not available to your local shell when you use tdownload; so your local shell cannot expand the wild cards.

trdate is a time setting utility. It reads the time on the remote machine and sets the local clock to the remote time. It must be run as root.

trdated is the daemon version of trdate. When it is started in ’rc.local’ it is run as daemon in which case it updates the time every 5 minutes (default). Even when there is no TERM connection, this daemon will start up when set in the rc.local. Once a TERM connection is created it starts updating the time.

6 X and TERM

TERM allows users to open up X windows on the local machine from clients that are running on a machine on the network. This is done by using the txconn client. txconn is executed on the remote, network-connected machine; it is invoked simply as txconn. It goes into the background and returns a number on the standard output; this number is the display number that clients should use to access the X server on the local machine. An example should make this clear. I am logged in, via trsh, to my remote TERM host, named foo. On foo, I do the following

  foo$ txconn
  Xconn bound to screen 10
  :10
  foo$

Now, on any host that I wish to run an X client on, that is to display on my local machine’s X server, I do

  setenv DISPLAY foo:10

(for bash you should use export DISPLAY=foo:10). In some cases it can furthermore be necessary to do a

  xhost + foo

7. tredir  23

or even a

  xhost +

on your local machine. Now when I start the client, it will try to connect to screen 10 on machine foo, but txconn is listening to this screen, and will forward all X protocol packets via TERM to the X server on the local host - i.e. the window will open up on your local machine.

It is possible to go the other way - run a client on your local machine and have it open up a window on a remote machine on the network; however we will defer explaining this until after we have discussed tredir.

The X protocol is not very efficient; it wastes some bandwidth. This is usually not a problem over an ethernet, but can be murder over a modem. X11R6 is supposed to introduce a low bandwidth version of the X protocol, LBX. If however you are using X11R5 you can use a utility named sxpc which compresses the X protocol, improving response over serial lines. Sxpc includes a write-up on how to get it working with TERM, and is recommended. The sxpc package also explains how to use xauth; so it is doubly recommended.

7 tredir

tredir is one of TERM’s most powerful utilities, allowing most important network services to be performed over a TERM link. Before we explain how to use tredir, it is necessary to give some background on network services. We have talked about network services before, but we have not said exactly what they are. Services are just that - services that are provided by the network. Examples of services include telnet, which provides logins between machines, the File Transfer Protocol, ftp, which transfers files between machines, and smtp, the Simple Mail Transfer Protocol, which is used whenever you send electronic mail. Each network service has a port number associated with it. The mapping of port numbers to services is given in the file ’/etc/services’. This file should be the same on all internet-connected machines.

How are these services invoked? Each networked machine runs a daemon called inetd, which listens for attempts to connect to the network ports. These requests can come from either the network or the local machine. A network service is obtained by connecting to a particular inetd port. When a network request is made, inetd knows exactly which service is involved by the port number the request is made on. If inetd is configured to do so, it provides the relevant service to the requesting connection. inetd’s configuration is given by the file ’/etc/inetd.conf’, which has a list of the services that inetd provides. For more information, see the man pages for inetd and inetd.conf.

You can communicate directly with network services by using telnet (n.b. not termtelnet). For example, to talk to the sendmail, (or smtp) daemon on machine machine name, you can do a telnet machine name smtp, or telnet machine name 25, (since 25 is the number assigned to smtp in ’/etc/services’). You should get a polite greeting from the daemon on the remote machine. This is a very useful trick for debugging network problems and checking ports redirected with tredir (see below).

Tredir works very much like inetd. It runs in the background as a daemon, listening to the network ports, waiting for a request. When a request for a service is made, instead of providing that service, as inetd does, tredir forwards the request over the TERM link to the remote TERM, which makes the request over the network, returning the result back over the link to the local client. Tredir can forward the request to any machine on the network, but by default sends it to the machine at the other end of the TERM link. Tredir redirects TCP (Transmission Control Protocol) network services over the TERM link.

The common command format of tredir is:

  tredir [this_computer:]port [that_computer:]port

An example should make this clear. Let us redirect a local port to the telnet port on the remote machine. To do this we would do

  tredir 2023 23

Now, anyone who connects to port 2023 on the local machine will be redirected to port 23 (telnet) on the remote machine. Here is an example session; the local machine is mymachine.modem.home and the remote machine is netsun.

  $ tredir 2023 23
  Redirecting 2023 to 23
  $ telnet localhost 2023
  Trying 127.0.0.1...
  Connected to mymachine.modem.home
  Escape character is '^]'.

  SunOS UNIX (netsun)
  login:

This example is actually quite useful. If I were instead to do the tredir on netsun. I could then telnet in to mymachine from the network simply by connecting to the redirected port on the networked machine (using

telnet) - i.e. telnet netsun 2023.

The general principle in using tredir is to redirect the desired service to a machine on the network. Our next example will allow us to read news on the local machine over our TERM link from a news server on the network. News is provided by the nntp service, port number 119. All decent news readers allow you to specify what port number they will use, either via a configuration file or an environment variable. Let us specify this local port to be 2119. Now, let us say that our news server is news.domain.org. We will redirect port 2119 to port 119 on news.domain.org; we will then tell our news reading software that the nntp server is located at port 2119 on the local host. Since this will depend on the news reader that you use, I will just test the link with telnet instead of firing up a news reader:

  $ tredir 2119 news.domain.org:119
  Redirecting 2119 to news.domain.org:119
  $ telnet localhost 2119
  Trying 127.0.0.1...

  Connected to mymachine.modem.home.
  Escape character is '^]'.
  200 news.domain.org InterNetNews NNRP server INN 1.4 07-Dec-41 ready
  (posting ok).

If you can get this far, all you have to do is configure your news reader to be able to read news via TERM. (n.b., if you read news like this, be sure that in all your posts you set a Reply-To: header to an network email address that you can be reached at, otherwise people who want to get in touch with you will be sending mail to whatever (wrong) data your news reader puts in the From: header).

7.1 tredir can bite!

The astute reader, after reading the last example will be wondering why port 2119 was redirected to port 119 - since news readers default to port 119, why could I not do a tredir 119 news.domain.org:119 and skip the news reader configuration? The answer is that all ports numbered less than 1024 are “reserved ports”, and only the superuser can listen to them. If one is willing to take a security risk and make tredir an SUID program, or run tredir as root, then one can redirect reserved ports and avoid the hassle of renaming services.

Another problem with using reserved ports is that inetd is often already listening to these ports, and only one program at a time can listen to a port. In order to use such a port, you must change ’inetd.conf’ so that inetd no longer listens to the port you want to redirect. This is most easily done by commenting out the line with the offending service by putting a # character at the beginning of the line. The superuser must then send inetd a HUP signal (kill -1 inetd-pid) to get it to reread its configuration.

7.2 Stupid tredir tricks

In this section we will describe some of the more common uses for tredir. We have already described how to redirect nntp and telnet services; here we will give some more complicated examples.

7.2.1 X windows

In a previous section, we described how to get an X client running on the network to open a window on your home machine using txconn. The same technique could be used on your home machine to display a client on the machine at the remote end of your TERM link. But how does one display an X client on a network machine that is not the remote end? The answer lies in knowing that X uses a particular network service just like the other programs we have been discussing. An X server listens for a network request on a port whose number is given by the formula port = 6000 + displaynumber, e.g. an X server managing screen 0 on a machine would listen to port 6000, if it were managing screen 2, it would listen to port 6002. When you set your DISPLAY environment variable to xmachine:n, your X clients will try to connect to port 6000 + n

on xmachine.

We can use this to trick X clients on your local machine to open up windows on remote displays. Let us say I want to open up an xterm, running on my local machine, on display 0 of machine xmachine, which is running some place on the network. I first pick a local display number, say 2 (do not use 0, since that is what your local X server will be using). I will map this display to display 0 on xmachine. In terms of ports, this means I want to redirect the local port 6002 to the remote port 6000. I do the following

  $ tredir 6002 xmachine:6000
  $ setenv DISPLAY localhost:2
  $ xterm

This should open up an xterm on machine xmachine. Note that I set the DISPLAY to localhost:2. This is because X clients will sometimes use unix domain sockets instead of internet domain sockets, at their own option, when connecting to a local display, if DISPLAY is set to :2. localhost:2 says to use a TCP connection.

Note that as far as xmachine is concerned, the X request is coming from the machine on the remote end of your TERM link (remotemachine) - so if you need to authorize the connection, you should either do an xhost + remotemachine on xmachine or use xauth to update the ’.Xauthority’ file on your local machine for display number 2, using the key from xmachine.

Again, to speed up X connections, you can use the program sxpc, which includes an explanation of how to use tredir to establish the link and authorize it using xauth.

7.2.2 Mail with TERM

Well, you asked for it. Electronic mail has the justifiable reputation of being one of the most difficult things to get working right on a UNIX system. To really get TERM working correctly with mail means that you have to understand how mail works, which is beyond the scope of this document. To learn more about mail, you should consult a book on UNIX system administration and/or the comp.mail.misc FAQ, available for anonymous ftp on rtfm.mit.edu:pub/usenet/comp.mail.misc. There are also currently two packages available for anonymous ftp on sunsite.unc.edu that will help you get mail running under TERM - they are

term.mailerd+smail by Byron A. Jeff and the BCRMailHandlerXXX by Bill C. Riemers.

That being said, we will give a thumbnail description of how mail works. There are two parts to getting mail running, sending messages and receiving messages. We will begin with sending messages from your local box to the network.

There are two classes of mail programs. The first is the mail user agent (MUA). MUA’s help you read, compose and send messages. Examples of MUA’s are elm, pine, Mail and vm. MUA’s do not really do any networking; they just put the messages together - the real work of sending mail is done by the second class of mail programs, the mail transfer agents (MTA’s). These are invoked by the MUA’s. They take the message, decide where to send it by looking at the address, and then actually deliver it over the network.

The two most common MTA’s on Linux systems are sendmail and smail. The basic idea is to get your MTA to connect to another MTA running on a machine on the net that will know what to do with your message. This is done by redirecting a local port to the smtp port on the net machine. You then have to tell you MTA to take any message it does not know what to do with, and send it out over the redirected port on your local machine to the MTA on the remote machine, which will then route your message to its correct destination.

How do we do this using smail? We first redirect a port to the smtp port on the network mail machine

(mailhost):
tredir XXXX mailhost:25

here XXXX is the port number that the smail on the localhost will connect to (note that I have to give this port a name in my ’/etc/services’ to get smail to recognize it). Smail has several configuration files that usually live in ’/usr/local/lib/smail’. The ones we care about are ’config’, ’routers’ and ’transports’. Note that I am assuming you have already got smail configured correctly for local mail - delivery to files and pipes and such things. Again, consult the documentation if you have not.

In the file ’config’, we put the following definition:

  smart_path=localhost

localhost is the machine that smail connects to when it does not know what to do with a message.

In ’routers’ we put

  smart_host:
  driver=smarthost,
  transport=termsmtp;
  path = localhost

In ’transports’ we put

  termsmtp:        driver=tcpsmtp,
        inet,
        return_path,
        remove_header="From",
        append_header="From: YOUR_NET_ADDRESS",
        -received,
        -max_addrs, -max_chars;
        service=YOUR_SMTP_SERVICE,

In the above, the header lines change the From header in all your outgoing mail to the address, YOUR NET ADDRESS, which is the network address you want mail sent to. If more than one user is going to be using your TERM link, you will have to do something more fancy, like keep a database of local user’s network addresses and insert these in the From: headers.

The service line is the name of the local port number that you have redirected to the smtp port on the network connected machine. In my version of smail, I cannot just set this to a number, I have to set it to a name, like “foo”, and then define “foo” in my ’/etc/services’ to be the number of my redirected port. If you use a SUID tredir and just redirect the smtp port (25), you do not need to define this.

This should be enough to get you going. If you decide to use sendmail, the principles are the same but the details differ. Ronald Florence (ron@mlfarm.com) told me that the stock Sun sendmail will not send multiple queued messages over a redirected port; BSD sendmail 8.6.9 works fine. He made the following changes to ’/etc/sendmail.cf’ to get it working with TERM. In his case, the default sendmail port (25) is used for SMTP traffic over a local ethernet so Internet mail is forwarded to a redirected TCP port.

  #
  #Create the termsmtp mailer, which sends mail via a re-directed TCP port
  #
  Mtermsmtp,P=[TCP], F=mDFMuCXe, S=22, R=22, A=TCP $h PORTNUMBER

Here, PORTNUMBER is the number of the redirected port on the local machine. This should be an unused port over 2000. We next tell sendmail which machine to connect to, and set termsmtp as the default mailer.

  #
  # major relay mailer
  #
  DMtermsmtp
  #
  # major relay host: use the $M mailer to send mail to other domains
  #
  DR HOSTNAME
  CR HOSTNAME

Here HOSTNAME is the name of your local host (does localhost work?). The last entry goes under Rule 0 to forward Internet mail.

  # Pass other valid names up the ladder to our forwarder
  R$*<@$*.$+>$*         $#$M    $@$R $:$1<@$2.$3>$4     user@any.domain

When the TERM connection is established to the Internet host, run the following commands on the local machine.

  tredir PORTNUMBER internet.host:25
  /usr/lib/sendmail -q

We now turn to receiving electronic mail using TERM. We will assume that mail is sent to your account on the network machine mailhost. The simplest solution is to just use trsh or termtelnet to log on to mailhost and read your mail on there. However, it is also possible to have your mail automatically downloaded to your local machine. One way to do this is to use the Post Office Protocol, (POP). POP was designed for exactly this purpose: to deliver mail to machines that have intermittent network connections. To use POP, you must have a POP server installed on mailhost. Assuming that you do, you can then use a POP client to download your mail every few minutes. This is done, as you might expect, using tredir. The POP service is 110 (note that there is an older protocol, POP-2, which uses port 109; in this document we describe

8. tudpredir  29

POP-3, which is the latest version of POP). There are several POP clients available. One, written in the script language perl is pop-perl-1.X, written by William Perry and maintained by myself - it can be found on sunsite.unc.edu:pub/Linux/system/Mail.

To use POP, you redirect a local port to port 110 on mailhost and configure your client to retrieve your mail from localhost using the local port. As an example, we will assume that there is a POP server running on mailhost. We will redirect the local port 2110, and fire up the pop-perl client:

  $ tredir 2110 mailhost:110
  Redirecting 2110 to mailhost:110
  $ pop
  Username: bill
  Password: <enter your password for mailhost>
  Pop Host: name of local
  Pop Port: 2110
  Starting popmail daemon for bill

If you do not have a POP server available, the BCRMailHandler package has a program to download your mail over a TERM link to your local machine. I have not used it, but anyone who has is welcome to comment. You can also use the term.mailerd+smail package for this purpose. Alas, both BCRMailHandler and the term.mailerd.smail package do not work anymore with TERM versions 2.0.0 and higher.

8 tudpredir

tudpredir is similar to tredir when you look at what these programs do and how they are executed. The big difference between the two programs is that tredir is used to redirect TCP network services while tudpredir redirects UDP (User Datagram Protocol) network services over the TERM link. One more important difference between the two programs is that tredir becomes a background daemon once it has successfully established the local port, while tudpredir commands must be placed in the background manually.

The common command format of tudpredir is:

  tudpredir [this_computer:]port [that_computer:]port

9 Automating Things

Now that you know how to get all your network services over TERM, it would be nice to set things up in such a way that your link is set up and configured automatically. There are basically an infinite number of ways of doing so, depending on what communication program you use and how you log in to your remote system.

One program that I have not used, but I have heard is quite nice, is fet: a front end for TERM. It is designed to log you into a remote system and fire up TERM and all your tredir’s. Any comments on fet would be most welcome.

I shall now give an example of a set of commands that use kermit to log into the remote system and then performs all of the TERM initializations. Obviously, if you use these examples, you will have to modify them for your own login procedures.

The command which is actually invoked is the shell script ’knet’, given by:

#!/bin/sh
/usr/bin/kermit -y $HOME/.kerm_term > $HOME/klog < /dev/null 2>& 1
exec $HOME/bin/tstart >> $HOME/klog 2>& 1

The script ’.kerm term’ is given by:

pause 2
# The number you want to dial
output atdtXXXXXXX \13
# Login to the terminal server
input 145 {name: }
output MYNAME \13
input 3 {word: }
output MYPASSWORD \13
input 5 {xyplex>}
# Make the line transparent
output term telnet-t \13
output term stopb 1 \13
# Connect to the remote host
output telnet remotehost.somedomain.org \13
input 10 {ogin: }
output MYOTHERNAME \13
input 3 word:
output MYOTHERPASSWORD \13
pause 5
# Fire up term on the remote host
output exec term -s 38400 -l $HOME/tlog -w 10 -t 150  \13
! /usr/bin/term -r -l $HOME/tlog -s 38400 -c off -w 10 -t 150 < /dev/modem  > /dev/modem &
# Open other clients here
suspend
!killall -KILL term

and finally, the script ’tstart’ which fires up the TERM clients is given by

#!/bin/sh
#
# This lets mail get out, can read news here, can pick up my mail here

#
/usr/local/bin/tredir 2025 25 2119 newshost:119 2110 pophost:110
#
# So I can open up Xwindows here
#
/usr/local/bin/trsh -s txconn
#
# So I will receive mail....
#
/usr/local/bin/pop
#
# Clean out the queue, in case of boo-boos
#
/usr/bin/runq
#
# Done now
#
echo ^G^G > /dev/console

When finally you want to close the connection, you resume and terminate kermit. The last line of the script kills the local TERM and returns the system in its initial state.

(Note of the author:  instead of doing '!killall -KILL term', I think it should be
possible just to do '!tshutdown'.  This should also work?)

As I said, there are zillions of ways to do so; these are just meant as examples to get you started. Other examples can be found in the packages autoterm and JoelTermStuff.

10 Porting software for use with TERM

In principle, all programs that can be used over a network can also be used in combination with TERM. Some of them you can get as binaries with TERM support already built in. These include telnet, (nc)ftp, Mosaic and many others. Most of these programs are compiled for TERM 1.17 or earlier. They should, however, still work with the newer versions of TERM.

Another way to make programs TERM aware is to port them yourself. This process will be described in the next subsection.

The last way to make your programs TERM-aware is to termify them.

10.1 Port and compile the sources

Porting software to TERM can be done using a fairly simple porting procedure:

If installed in ’/usr/local’ by root:

  1. Add to the compile flags -include /usr/local/include/termnet.h

2. and add to the library list -ltermnet

If installed in your home directory:

  1. Add to the compile flags -include $HOME/term/termnet.h

2. and add to the library list -L$HOME/term -ltermnet

Now compile the software as described in the INSTALL or README document that came with the software. That should do!

At this point the commands should work both with and without TERM.

  telnet localhost

does not use TERM to connect, but

  telnet bohr.physics.purdue.edu

will use TERM only if you do not have some other type of network connection.

Some commands like rlogin can only be executed by root and the owner of the TERM connection (privileged persons).

Some TERM commands will be TERM transparent and only use TERM when there is not another option. Some common examples are telnet and ftp.

Others require an external flag to tell them it is all right to use TERM. These programs include xarchie, fsp and ytalk.

You can either flag these programs to use TERM by setting the environmental variable TERMMODE as specified in README.security, or running make installnet. Eventually, the ’termnet’ file created will contain special networking instructions, but for now only its existence is checked.

If you add an ethernet connection, you can then simply remove the ’termnet’ file and continue to use the same binaries!

NOTE: Programs that were ported back in the days of client.a, can still be recompiled for use with newer versions of TERM simply by changing the client.a reference to libtermnet.a.

10.2 Termify

This package will convert dynamically linked binaries for TERM use.

Before you can use termify you have to make sure that you have TERM version 2.2i (is this version 2.2.8?) or later and libc.so.4.5.26 or later. Then you have to create the file ’libt.so.4’ in the directory ’/lib’ (see the README file in the package).

Problem at this moment is that you have to remake the file ’libt.so.4’ every time you upgrade TERM versions.

After you have created the library you can let termify ’crunch’ the program you want to make TERM-aware, by using the command:

  termify <command name>

When you do not like the result you can ’un’termify the program you have just termified using the command:

  termify -u <command name>

Last, the package also contains a script for completely termifying ’smail’; so no special transport definitions are necessary. The only thing you possibly want to change is the ’From: ’ address.

11 Term clients

11.1 Term clients available on ftp sites.

Below a list of application running with TERM is given. I am not stating that this list is complete; so any completion is welcome. As far as possible I will present the site and directory where the application can be found (all to my knowledge). When I state sunsite.unc.edu as the place to find the application I mean that you can find it in one of the following two directories:

1. /pub/Linux/apps/comm/term/apps

2. /pub/Linux/apps/comm/term/extra

Here we go :-)

TERM package:

     tupload
     tdownload           (versions 2.1.0 and higher)
     trsh
     tmon
     tredir
     tudpredir           (versions 2.0.0 and higher)
     txconn
     trdate(d)
     tshutdown
     libtermnet

File transfer:

11. Term clients 34

ftpd                sunsite.unc.edu
termncftp           sunsite.unc.edu
ncftp185            sunsite.unc.edu:/pub/Linux/system/Network/file-transfer
fsp                 sunsite.unc.edu:/pub/Linux/system/Network/file-transfer

Information systems:

lynx
Mosaic              sunsite.unc.edu:/pub/Linux/system/Network/info-systems/Mosaic
chimera
netscape            sunsite.unc.edu:/pub/Linux/system/Network/info-systems
httpd
xgopher
gopher              sunsite.unc.edu

Remote login:

termtelnet          sunsite.unc.edu
rlogin              physics.purdue.edu:/pub/bcr/term/extra
rsh                 physics.purdue.edu:/pub/bcr/term/extra

Netnews:

tin 1.3             sunsite.unc.edu:/pub/Linux/system/Mail/news
news2               sunsite.unc.edu

Mail:

slurp               sunsite.unc.edu
smail               sunsite.unc.edu
term.mailerd+smail  sunsite.unc.edu
BCRMailHandlerXXX   physics.purdue.edu:/pub/bcr/term

Automating scripts:

JoelTermStuff       sunsite.unc.edu
autoterm            sunsite.unc.edu
fet                 sunsite.unc.edu

Other programs:

inetd               sunsite.unc.edu
rdate               sunsite.unc.edu
xgospel             sunsite.unc.edu:/pub/Linux/games/x11/networked
termify             physics.purdue.edu:/pub/bcr/term/extra
xboard              sunsite.unc.edu

     ircII               sunsite.unc.edu:/pub/Linux/system/Network/chat
     whois
     xwebster            sunsite.unc.edu
     sxpc                ftp.x.org:/R5contrib
     xztalk              sunsite.unc.edu:/pub/Linux/apps/sound/talk

11.2 The termnet package

The package termnet-2.0.4-Linux-bin.tar.gz (sunsite.unc.edu:/pub/Linux/apps/comm/term) contains a couple of pre-compiled TERM clients and a couple of scripts, manual pages and libtermnet.so.2.00.04. The clients are compiled using this version of libtermnet.so. The package contains the following clients:

     fet         perl        sperl4.036  tmon        tshutdown   xgopher
     finger      perl4.036   suidperl    trdate      tudpredir   ytalk
     ftp         rcp         taintperl   trdated     tupload
     fwhois      rlogin      telnet      tredir      txconn
     ncftp       rsh         term        trsh        xarchie

WARNING: The package also contains the complete set of compiled clients of TERM 2.0.4 including TERM itself. Do not install this package before you are sure about what you want. You will destroy other versions of TERM and its clients when you start moving executables around.

11.3 Asked for but not yet supported

1. DOOM: The problem with this game seems to be the fact that it uses port 5029 both as client and as server.

2. NFS: The NFS server is only supposed to accept requests if the socket requesting the connection is bound to a port below 1024. This seems to be troublesome. However, some NFS servers have an ’insecure’ option. In this case NFS might work eventually, if RPC support is added to Term.

12 Term and Security

In this section I will point to some security aspects of TERM. The problems will be explained and a way to improve security will be given.

12.1 trsh

Trsh is insecure when it is used to access the local Linux box from the remote system. The problem with TERM and its clients is that beside the owner of the TERM connection also ’root’ can execute TERM-aware programs over the connection.

This also means that ’root’ on the remote system can execute trsh and thus can enter the login which owns the TERM connection quite easy. If this owner on the local box is ’root’ then you will be in hell.

The solution to this problem is easy: you just have to put the following line in the termrc file on the local box:

  denyrsh on

With this set in the ’termrc’ file, nobody can use trsh on the remote site anymore to access your machine. When you and others want to access your local Linux box over the TERM connection this can still be done using telnet and redirected ports.

12.2 txconn and xauth

Txconn is not terribly secure; anyone can connect to your local server via TERM and perform all sorts of mischief. I you are worried about this sort of thing, it might be a good idea to consider using xauth to authorize your connections. See the next section for an example of using xauth for securing your connections.

12.3 sxpc, xhost and xauth

Sxpc in combination with ’xhost +’ is very dangerous when you are not using xauth.

Using xauth is very important to maintaining security when using sxpc. If you do not use xauth when using sxpc all the dangers of running with ’xhost +’ apply. These dangers include but are not limited to:

• Someone watching what is displayed on your screen

• Someone watching what you type

• Someone typing in one of your windows (for example: a command to delete all your files :-(

Xauth is available in X releases R4 and later. Here I will describe how to set up basic usage of xauth. This configuration is vulnerable to network snooping, but if you can live with that it should be fine.

NOTE: when using xauth your $DISPLAY variable must NOT be set to localhost (or localhost:whatever). If your $DISPLAY variable does use localhost the clients will be unable to find the appropriate authorization information. The workaround is to use the real hostname of the machine. If you follow the compilation instructions in the README, and compile without -DNOGETHOSTNAME then everything should work.

The machine where you will be running clients will be called C, the machine where you wish to display them will be called D.

First choose a ’key’, up to 16 pairs of hexadecimal digits (so an even number of characters from the ranges 0-9 and a-f). You will need to supply this key in place of <key> in the example below.

On C:

  % xauth
  xauth:  creating new authority file $HOME/.Xauthority
  Using authority file $HOME/.Xauthority

  xauth> add Chostname:8 MIT-MAGIC-COOKIE-1 <key>
  xauth> exit

On D:

  % xauth
  xauth:  creating new authority file $HOME/.Xauthority
  Using authority file $HOME/.Xauthority
  xauth> add Dhostname/unix:0 MIT-MAGIC-COOKIE-1 <key>
  xauth> add Dhostname:0 MIT-MAGIC-COOKIE-1 <key>
  xauth> exit

When starting the X server on D you should give the flag -auth $HOME/.Xauthority. You may need to edit or create a ’$HOME/.xserverrc’ to control how the X server is started. For example:

  #!/bin/sh
  exec X  -auth $HOME/.Xauthority $*

Make sure that your ’.Xauthority’ file is readable only by you on both C and D.

13 Things to remember

In this section I try to present you with a list of useful ftp addresses, URL’s etc. where you can find software and information about TERM.

Ftp:

sunsite.unc.edu:/pub/Linux/apps/comm/term/<whole-directory-tree>
sunsite.unc.edu:/pub/Linux/docs/HOWTO/<whole-directory-tree>
physics.purdue.edu:/pub/bcr/term/<whole-directory-tree>

URL:

http://sunsite.unc.edu/mdw/HOWTO/Term-HOWTO.html
http://www.bart.nl/~patrickr/term-howto/Term-HOWTO.html  (always the newest version)
http://physics.purdue.edu/~bcr/homepage.html

Netnews:

comp.os.linux.announce           announce of new TERM versions and Term-HOWTO
comp.os.linux.help               ask your questions about TERM here
comp.os.linux.misc               or here
comp.protocols.misc              answers to TERM questions are also posted here.

Related documents:

Using Term to Pierce an Internet Firewall HOWTO    by Barak Pearlmutter,
                                                      bap@learning.scr.siemens.com
Firewall HOWTO                                     by David Rudder,
                                                      drig@execpc.com
Serial HOWTO                                       by Greg Hankins,
                                                      gregh@cc.gatech.edu
Net-2/Net-3 HOWTO                                  by Terry Dawson,
                                                      terryd@extro.ucc.su.oz.au

When you start asking questions on netnews please make sure that you give people in the groups as much information as they need to solve your problem (TERM version number, way you set up your connection etc.). At this moment many TERM versions are in use and all have their specific and common problems. Therefore, when you want a useful answer, at least state the version of TERM you are using. Otherwise, in some cases only wild guesses are possible to help you solve your problems.

14 Reliability of TERM versions

Many versions of TERM are around now. The maintainer of TERM, Bill Riemers, has made a list of TERM versions stating what versions are reliable and what versions you had better avoid. This list is the following:

             term110          --> cannot really say
             term111          --> cannot really say
             term112          --> cannot really say
             term113          --> cannot really say
             term114          --> fairly stable BETA version
             term115          --> unstable BETA version
             term116          --> unstable BETA version
             term117          --> unstable BETA version
             term118          --> semi-stable BETA version
             term119          --> stable GAMMA version
             term-2.0.X       --> semi-stable BETA versions
             term-2.1.X       --> more stable BETA versions
             term-2.2.X       --> new BETA versions
             term-2.3.X       -->

15 Term speed table

Thanks to Bill McCarthy we now have a table with TERM speed information for different modems, TERM versions and connection conditions. Its purpose is to give new and experienced users some idea what other people are using and the results they are getting.

15. Term speed table 39

                           LINUX TERM CHART 8/14/94

|___modem speed/make___|___line speed__|__avg cps__|__high__|__term ver_|
| 1)  USR SP 14.4      |   9600        |  950      |  963   | 1.17      |
| 2)  USR SP 14.4      |  14400        | 1376      |  n/a   | 1.18p06   |
| 3)  Zoom 2400        |   2400        |  220      |  230   | 1.19      |
| 4)  Boca V.32bis 14  |  57600        | 1400      |  n/a   | 1.01/09?  |
| 5)  Viva 14.4        |  14400        | 1300      |  n/a   | 1.16      |
| 6)  USR SP 14.4      |  14400+       | 1550      | 1680   | 1.19      |
| 7)  Intel 14.4 Fax   |  14400        | 1400      | 1650   | 2.0.4     |
| 8)  cable tv hookup  |  57600        | 1500      | 1800   | 1.18p06   |
| 9)  Twincom 144/DFi  |  57600        | 1500      | 4000?  | 2.0.4     |
| 10) USR SP 14.4      |  14400        | 1200      | 1500   | 1.08      |
| 11) cable tv hookup  |  19200        | 1300      | 1800   | 1.19      |
|-----------------------------------------------------------------------|
+Command flags/termrc settings:
1) default escapes   2) window 5       3) baudrate 2400    4) n/a
   baudrate 9600        timeout 200       window 3
   window 10                              noise on
   timeout 150
5) compress off      6) baudrate 19200  7) ignore 19+17    8) compress off
   window 10            compress on        window 4           escape 0, 13,
   timeout 150                             timeout 90         16-19, 255
   baudrate 38400                                             baudrate 0
                                                              shift 224
                                                              flowcrtl 500
                                                              window 10
                                                              timeout 70
                                                              retrain on
                                                              breakout 24
9) compress off     10) compress off      11) baudrate 19200
   baudrate 57600       baudrate 38400        compress on
   window 10            escape 17, 19         shift 224
   timeout 200          remote                escape 0, 13 16-17
   noise on                                   19, 255
   share on                                   window 10
   remote                                     timeout 40
Escaping characters at one end also implies ignoring them on the other end.

16 Hints and Tricks found on the net

In the Linux related newsgroups many questions about TERM are coming back every couple of weeks, together with the answers to these questions. To reduce traffic to the newsgroup, in this section I shall try to make a composition of these questions and the answers to them. Some of the answers have been checked by me as I also had related problems. Others, I have just taken from the newsgroups without testing them.

• Many people, especially those who are using Ultrix, seem to have trouble with vi presenting less than 24 lines on a window with 24 lines. There are three ways to get rid of this problem:

1. Log in into the remote system using:

      trsh -s telnet {\(<\)}hostname>

2. Put ’resize; clear’ in your ’.login’ file

3. The best solutions seems to be to enter the following remote:

      stty 38400

• Many people seem to have problems with crashing TERM connections, whatever may be the reason of the crashes. So before starting applications people want to know whether their TERM connection is still alive or not. This can be checked using the following small shell script examples: When you are using tcsh:

  if ( { trsh -s true } ) then
    ...
  endif

When you are using bash:

  if trsh -s true; then
    ...
  fi

• The WWW browser Netscape is causing people problems to get it to work with TERM. The good news is that it will work under TERM. Here is how:

        1. Termify netscape
        2. Fire up termnetscape
              under Options | Preferences | Mail/Proxys
              leave _all_of the proxy boxes blank
              set the SOCKS box to 'remotehost' & 80
        3. Ignore the error you get when you exit the Options menu.
        4. If termnetscape fails to work right:
              under Options | Preferences | Mail/Proxys
              leave _all_of the proxy boxes blank
              set the SOCKS box to 'none' & 80
        5. Ignore the error you get when you exit the Options menu.

Scott Blachowicz mailed me stating that it is even easier to get Netscape or any other browser to work with TERM if you have a proxy server (e.g. CERN’s httpd) installed for use on the remote side of your TERM link. On the local side, you can then do this:

             1. tredir localhost:8080 remotehost:80
                   where 'remotehost' is the hostname of your proxy server system.
             2. Run your particular browser, then in the preferences or options or
                   whatever it is called, point your proxies at 'localhost' port
                   8080.
                   Some browsers might like environment variables like
                       export http_proxy=http://localhost:8080/
                       export ftp_proxy=http://localhost:8080/

17 Other Things

Some things that might be included:

• Extension of troubleshooting

• Extension of security issues

• Termwrap

• Suggestions

Anyway, if you have suggestions, criticism, suggestions, or anything else to say about this document, please fire away. At the moment I, Patrick Reijnen, have taken over the authorship of the TERM-HOWTO. I can (currently) be reached at patrickr@bart.nl.

18 Acknowledgements

A lot of people have to be thanked. First and foremost Michael O’Reilly and all the developers of TERM, who have provided us with such a great tool. I would also like to thank everyone who gave feedback and contributed to this HOWTO. They include Bill Reynolds, the former author of this HOWTO, Ronald Florence, Tom Payerle, Bill C. Riemers, Hugh Secker-Walker, Matt Welsh, Bill McCarthy, Sergio, Weyman Martin and everybody I forgot to mention.